OpenStudy (anonymous):

Use the Laws of logarithms to rewrite the expression log2(2x(x−16)) in a form with no logarithm of a product, quotient or power. After rewriting we will have: log2(2x(x−16))=log2^A+log2x+log2f(x) what is the constant (A) what is the function f(x)

OpenStudy (across):

\[\log_{2}(2x(x-16))\]¿Es esto con lo que necesitais ayuda?

OpenStudy (anonymous):

necesito A and B

OpenStudy (across):

So you want me to re-write that thing as follows:\[\log_{2}(A)+\log_{2}(2x)+\log_{2}(f(x))?\]

jimthompson5910 (jim_thompson5910):

\[\Large \log_{2}\left(2x(x-16) \right)\] \[\Large \log_{2}\left( 2\right) + \log_{2}\left( x\right) + \log_{2}\left( x-16\right)\] So A = 2 and f(x) = x-16